A standing wave with wavelength λ = 1.2 m and frequency f = 50 Hz is generated on a stretched cord. For an element of the cord at x = 0.5 m, the maximum transverse velocity is v(y,max) = 2π m/s. The amplitude A of each of the individual waves producing the standing wave is:


0.0125 m


0.03 m


0.02 m


0.01 m


0.025 m


1
Expert's answer
2021-05-26T10:37:20-0400

Gives

"\lambda=1.2m"

"f=50Hz"

"x=0.5m"

"V_{max}=2\pi" m/sec

We know that

Individual wave equation

"y=Asin(wt-kx)\rightarrow(1)"

eqution differenciate with respect time

"V=Awcos(wt-kx)"

Phase "(wt-kx)=0\u00b0"

Then "cos(wt-kx)=cos0\u00b0=1"

"V=Aw\rightarrow(2)"

"2\pi=Aw"

"2\pi=A2\pi\times50"

"A=\frac{1}{50}=0.02m"

Amplitude of each standing wave(A)=0.02m

Option (c) is correct option